OpenStudy (anonymous):

Simplify: 12(5c - 3).

OpenStudy (anonymous):

\[12(5c-3)\] apply distributive property a(b+c) =ab+ac hence \[(12)(5c) - (12)(3) = 60c - 36\]

OpenStudy (anonymous):

NORAS how can u find the value of c with no equal to sign given i mean it is an expression not an equation

OpenStudy (anonymous):

C=0.6

OpenStudy (anonymous):

HEHEHE

OpenStudy (anonymous):

oh come on noras it may confuse the math master :)

OpenStudy (anonymous):

\[60C=36 \]

OpenStudy (anonymous):

what noras 12(5c-3) is given no equal to sign ...

OpenStudy (anonymous):

\[C=36\div60=0.6\]

OpenStudy (anonymous):

I GIVE THE EQUAL BY MY SELF

OpenStudy (anonymous):

HEHEHE JUST KIDDING ANYWAY

OpenStudy (anonymous):

heheheh